3sqrt x2=9
answer:
Show work:

Answer by stanbon(48535)

3sqrt x^2=9
Divide both sides by 3 to get:
sqrt x^2=3
Square both sides to get:
x^2=9
Solve for x to get:
x=3 or x=-3
========
Cheers,
Stan H.